OpenStudy (lulubj):

The vertex of the parabola is at (-4,-1). When the y-value is 0, the x-value is 2. What is the coefficient of the squared term in the parabola's equation? A. 6 B.-6 C.-3 D.3

OpenStudy (anonymous):

The vertex equation for a parabola with vertex (h, k) is \[y = a(x-h)^2+k\]

OpenStudy (anonymous):

Using the given vertex (-4, -1), \[y = a(x+4)^2-1\]

OpenStudy (anonymous):

Do you agree?

OpenStudy (anonymous):

Plug in (2, 0) into the parabola equation to solve for a.

OpenStudy (lulubj):

Thanks

OpenStudy (lulubj):

So would it be 6?

OpenStudy (anonymous):

\[0 = a(2+4)^2-1 = 36a - 1\]

OpenStudy (anonymous):

\[36a = 1\]